I collect a sample from a population and I want to make a confidence interval around the sample mean. The sample variance is 16, and my sample size is 16. If I want to make a 95% confidence interval, what would my margin of error be? (hint: margin of error is just the part after the +- in the formulas)
Answer #1

We have given,

Sample variance = 16

Sample size =n=16

Degree of freedom =n-1=16-1=15

t critical value = 2.131...............by using t table or Excel =TINV(0.05,15)

Margin of error

  

=2.131